OpenStudy (anonymous):

A bag contains 7 blue counters, 5 green counters, and 3 pink counters. The probability of randomly drawing a green counter is 1/3. What is the probability of not drawing a green counter

OpenStudy (anonymous):

help please :0

OpenStudy (anonymous):

probability of not drawing a green counter = probability of drawing blue counters + pink counters

OpenStudy (anonymous):

or 1 - probability of drawing a green counter

OpenStudy (anonymous):

2/3 is the answer

OpenStudy (anonymous):

so 1 is actually the sum of all probability (in case u need to write it down .... probability of blue + prob. of pink + prob. of green)

OpenStudy (anonymous):

\[probability \left( pink \right)+prob.\left( blue \right)+prob.\left( green \right)=your answer\]